Kimiya has 4times narbles as ramneet.if kimiya gave 18 to ramneet they wud have the same number. How many marbles has each?
Answer by TimothyLamb(4379)

x = kims number
y = rams number
---
x = 4y
x - 18 = y + 18
---
put the system of linear equations into standard form
---
x - 4y = 0
x - y = 36
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = kims number = 48
y = rams number = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php